# Mean-field theory of the DNLS equation at positive and negative absolute temperatures

### v1

@@ Abstract
The Discrete Non Linear Schrödinger (DNLS) model, due to the existence of two conserved quantities, displays an equilibrium transition between a homogeneous phase at positive absolute temperature and a localized phase at negative absolute temperature. Here, we provide a mean-field theory of DNLS and show that this approximation is semi-quantitatively correct in the whole phase diagram, becoming exact in proximity of the transition. Our mean-field theory shows that the passage from stable positive-temperature to metastable negative-temperature states is smooth.
keywords:
Negative temperature, mean-field theory, phase transitions
@@ 1 Introduction
The Discrete Non Linear Schrödinger (DNLS) equation is around fifty years old, but it continues to attract the interest of researchers [1, 2, 3] for at least two reasons. On the one hand, it has applications in a variety of different fields, ranging from the foundations of statistical mechanics [4] to solid-state physics [5]. On the other hand, its nonlinear and non integrable character is the basis of interesting non-trivial properties: from a dynamical point of view, the existence of localized breathers [6], and from a statistical point of view, the presence of a negative-temperature ($T$) regime [7] in a large part of its phase diagram. The two properties are strictly related, because the negativeT-T phase is characterized by the spontaneous emergence of breathers, which are in principle metastable [8] (the “equilibrium” state corresponds to a single breather sitting on top of an infinite-temperature background [9]). However, because of the extremely slow breather dynamics [10] (due to the spontaneous appearance of an adiabatic invariant [11]), the asymptotic state may be effectively unachievable [9].
@@
According to Machlup’s criterion, the existence of a negativeT-T phase is a general property of statistical systems displaying a finite energy density at infinite temperature [12]. In the DNLS case, this occurs because of the existence of two conserved quantities bounded from below: the energy and the mass (or norm). The purely stochastic C2C model [7, 13, 14] (where the energy is locally the square of the mass) is an even simpler example and represents a very good training ground, since the absence of interactions between neighbouring sites makes it possible to derive some rigorous results [15, 16]. Furthermore, it helped establishing a connection with a wider class of stochastic models [17, 18], the so called Zero Range Processes [19].
@@
In this paper we introduce a mean-field (MF) description of the DNLS. The approximation is based on a tweak of the interaction term: we replace the product of variables $x_{n}x_{n+1}$ on neighbouring sites, with the product between the variable on a given site with its statistical average along the system, $x_{n}\langle x_{n}\rangle$. As a result, the grandcanonical partition function can be factorized, thereby obtaining explicit formulas and reaching a better comprehension of some properties of the DNLS model. The MF approximation is exact in proximity of the transition line separating the homogeneous ($T>0$) phase from the localized ($T<0$) one. The reason is that spatial correlations are absent and the average contribution of the interaction energy vanishes along such a curve. In addition, the MF approximation gives a consistent and semiquantitatively correct description in the whole phase diagram, including the $T=0$ curve and the (metastable) negative $T$ region. In Sec. 2 we present the model and the method. The main results of the MF approximation are discussed making a distinction between positive (Sec. 3) and negative (Sec. 4) temperatures, because in the latter case it is necessary to introduce a cutoff in the mass integral which defines the partition function. All the details of the calculation and of the simulations are deferred to the appendices.
@@ 2 The Model and its Mean Field Theory
We start by defining the DNLS model and recalling the most important properties of its phase diagram. As usual, we consider a one dimensional lattice of size $N$ and associate a local mass $c_{n}$ and a phase $\phi_{n}$ to each site $n$. The system Hamiltonian is
@@
In Fig. 1, we plot the numerical values of $h_{nl}$ for the DNLS for different temperatures as a function of the mass density $a$ and for $J=1$. Interestingly, they align almost perfectly along the corresponding isothermal lines of the C2C ($J=0$), showing that given the mass density and the temperature, the nonlinear energy $h_{nl}$ (approximately) coincides in the two models. This implies that the angular degrees of freedom contribute “simply” as an additional interaction energy $h_{int}$, bounded between 2aJ-2aJ in the ground state and 0 on the critical line.
Fig. 1: Value of $h_{nl}$ along three isothermal curves ($\beta=0.01,0.1,10$ from top to bottom) comparing the results of equilibrium numerical simulations (see A) of the DNLS equation (symbols) with the prediction of the MF theory (dashed lines) and with the energy of the C2C model (dotted lines) as derived in [21, 22].
Within the grand canonical formalism, the equilibrium properties of DNLSE can be extracted from the partition function
@@
Fig. 2: Relative difference between the MF partition function $z$ computed by numerically integrating the exact expression (9) and its analytical approximation $z_{k}$ obtained as a series expansion at order $k$ in $w$. The various symbols correspond to different $w$ values as from the legend. The value of $\beta$ is always equal to $\pm 0.4$ except for the lowermost curve, where $\beta=-0.01$. Notice that the sign of $\beta$ coincides with the sign of $w$. For negative $\beta$, the numerical integration is performed by imposing the same cutoff $c^{*}$ as in the C2C model.
@@ 3 Positive temperatures
Fig. 3: Comparison between MF theory (curves) and exact results obtained via grandcanonical simulations (symbols). Each curve corresponds to a different $T=1/\beta$, increasing from bottom to top according to the legenda. The size of the system is $N=100$ and details about simulations are given in A.
@@
Fig. 4: Same as in Fig. 3, but we now compare MF and exact numerical results in the planes $(\mu,a)$ and $(\mu,h)$ rather than in $(a,h)$. In this way a horizontal shift at low temperature is evident, due to a wrong determination of the chemical potential in the MF theory. Symbols: exact results from grandcanonical simulations. Full lines: MF approximation. Dashed lines: MF approximation corrected for the chemical potential: $\mu\to\mu-1$.
@@
Fig. 5: The contributions $h_{int}$ and $h_{nl}$ are presented versus the total energy $h$ suitably rescaled (see the main text). Symbols are the exact results from microcanonical simulations, as described in Appendix A, but with the removal of the interaction with the heat bath and starting from an initial condition with the chosen values of $a$ and $h$.
It is worth stressing that the same linear behaviors are expected in the MF approximation, because the limiting values of the two contributions, $y_{nl,int}(0,1)$, are the same, see B and C, and because the exact and MF theories compare very well in the $(a,h)$ plane, see Figs. 1,3. Nevertheless, it is not feasible to derive analytically the linear behavior of $y_{nl,int}(x)$ within the MF theory, because such theory works with the grandcanonical variables $\beta,\mu$ and their relation with $a$ and $h$ can be easily obtained numerically, but no closed-form solutions are available. However, we can use the expansions in $w$, which is what we are going to do now.
@@
Fig. 6: This figure intends to represent the relationship between the total energy $h$ of the system and its nonlinear part, $h_{nl}$. For small $\beta$ both quantities tend to $h_{c}=2a^{2}$, therefore we plot the relative differences between the two energies and the critical energy. The depedence of this ratio on $m$ is well accounted for by the MF result (dashed line).
@@
It is worth stressing that the approximation $R=4/(4+\pi m^{2}/2)$ is not valid for vanishing $T$. In this limit $m$ diverges, which would imply a vanishing $R$, while in reality $R$ goes to a constant,
@@ 4 Negative-temperature homogeneous states
Fig. 7: Sketch of the potential $U(c)$, Eq. (18), for $\theta=0$ and small, negative $\beta$.
The crossover from the first to the second regime occurs at $c_{1}\simeq\beta^{2}q^{2}/m^{2}$, while the crossover from the second to the third regime occurs at $c_{2}\simeq|m/\beta|$. For the potential to be partially confining, it is therefore necessary that $c_{2}$ is sufficiently larger than $c_{1}$. This implies $|m|\gg|\beta|q^{2/3}$, which, in the limit of small $\beta$, leads to $a\ll 1/|\beta|^{3/4}$ (see Eqs. (11,12) and a comment in the conclusions, next section). In this limit, $|U(c_{1})|\ll U(c_{2})$, so that the anisotropic term is fully negligible and the potential barrier matches the one derived in [20]. In this regime, one can set the cutoff mass equal to the potential maximum, $c^{*}=c_{2}$. However, we also see that for “large” $a$ (and fixed $\beta$), the potential is not locally confining and one must expect breathers to arise much more rapidly. This is at variance with the C2C model, where there is always a minimum in 0.
Fig. 8: MF potentials $U(c,\phi=0)=\beta c^{2}+2q\sqrt{c}-mc$, for different values of $\beta$ and $m$. The full (blue) line corresponds to $\beta=0$ and $U(c)$ is linear.
@@ 5 Conclusions
In this paper, we have developed a mean-field approach to derive explicit expressions for the most relevant observables describing the equilibrium state of the DNLS equation. The approximation consists in assuming that the interaction is gauged by the average amplitude of the neighbouring sites rather than by their actual value.
In a certain sense, the mean-field approximation is a step forward compared to the stochastic C2C model in which the coupling term is completely neglected. In fact, the MF theory is accurate not only in the small $\beta$ region, where the interaction energy is on average negligible: it turns out that the resulting approximation provide a good representation in a much wider region of the parameter space, including the correct determination of the $T=0$ curve,
It is also interesting to notice that the MF model provides an accurate description across the transition from positive to negative temperature, where instabilities are expected to arise. In fact, MF theories typically fail precisely in proximity of phase transitions. Here, the success of MF is due to the fact that the interaction energy is negligible in that area, and the transition itself occurs already in the absence interactions.
@@
Fig. 9: Comparison of exact numerical results (circles) and MF prediction (blue line) for negative $\beta=-0.01$ and different values of m=5,2.5,1.5,1m={-5,-2.5,-1.5,-1}, from left to right. Main: comparison in the plane $(a,h)$. Inset: comparison in the plane $(a,h-2a^{2})$. Here it is manifest we are in the negativeT-T region, because $(h-2a^{2})>0$. We also make evident the statistical error bar for the circles, obtained as spatial and temporal averages of grandcanonical simulations.
Fig. 10: As in Fig. 9, but now the comparison is done in the planes $(\mu,a)$ and $(\mu,h)$ rather than in the plane $(a,h)$.
Data availability
Reasonable requests for data will be fulfilled.
Declaration of competing interest
The authors declare that they have no known com- peting financial interests or personal relationships that could have appeared to influence the work reported in this paper.
CRediT authorship contribution statement
MG did most of numerical simulations and calculations. All authors have contributed equally to the planning, discussions, and writing.
Declaration of generative AI and AI-assisted technologies in the writing process
Authors have not used generative AI and AI-assisted technologies in the writing process.
@@ Acknowledgements
SI and PP acknowledge financial support from the Italian MUR PRIN2022 project “Breakdown of ergodicity in classical and quantum many-body systems” (BECQuMB) Grant No. 20222BHC9Z.

### v2

@@ Abstract
The Discrete Non Linear Schrödinger (DNLS) model, due to the existence of two conserved quantities, displays an equilibrium transition between a homogeneous phase at positive absolute temperature and a localized phase at negative absolute temperature. Here, we provide a mean-field theory of DNLS through a suitable approximation of the grandcanonical partition function which makes it factorizable. By comparing our mean-field results with numerically exact ones, we show that this approximation is semi-quantitatively correct in the whole grandcanonical phase diagram, becoming increasingly accurate in proximity of the transition line and exact along the line itself. Our mean-field theory suggests that the passage from stable positive-temperature to metastable negative-temperature states is smooth and allows for a more accurate description of the metastable region.
: Journal of Statistical Mechanics: theory and experiment
Keywords: Negative temperature; Mean-field theory; Phase transitions
@@ 1 Introduction
The Discrete Non Linear Schrödinger (DNLS) equation is around fifty years old, but it continues to attract the interest of researchers [1, 2, 3] for at least two reasons. On the one hand, it has applications in a variety of different fields, ranging from the foundations of statistical mechanics [4] to solid-state physics [5]. On the other hand, its nonlinear and non integrable character is the basis of interesting non-trivial properties: from a dynamical point of view, the existence of localized breathers [6], and from a statistical point of view, the presence of a negative-temperature ($T$) regime [7] in a large part of its phase diagram. The two properties are strictly related, because the negativeT-T phase is characterized by the spontaneous emergence of breathers, which are in principle metastable [8] (the “equilibrium" state corresponds to a single breather sitting on top of an infinite-temperature background [9]). However, the rising of breathers is an extremely slow process because for large and negative temperatures there is a high barrier preventing it [10] and once breathers appear, also their dynamics is extremely slow [11], due to the spontaneous appearance of an adiabatic invariant [12]. As a result, the asymptotic state may be effectively unachievable [9], making the study of the homogeneous, metastable state of special interest.
@@
According to Machlup’s criterion, the existence of a negativeT-T phase is a general property of statistical systems displaying a finite energy density at infinite temperature [13]. In the DNLS case, this occurs because of the existence of two conserved quantities bounded from below: the energy and the mass (or norm). The purely stochastic C2C model [7, 14, 15, 16] (where the energy is locally the square of the mass) is an even simpler example and represents a very good training ground, since the absence of interactions between neighbouring sites makes it possible to derive some rigorous results [17, 18]. Furthermore, it helped establishing a connection with a wider class of stochastic models [19, 20], the so called Zero Range Processes [21].
@@
In this paper we introduce a mean-field (MF) description of the DNLS. The approximation is based on a tweak of the interaction term: we replace the product of variables $x_{n}x_{n+1}$ on neighbouring sites, with the product between the variable on a given site with its statistical average along the system, $x_{n}\langle x_{n}\rangle$. As a result, the grandcanonical partition function can be factorized, thereby obtaining explicit formulas and reaching a better comprehension of some properties of the DNLS model. The MF approximation is asymptotically exact in proximity of the transition line separating the homogeneous ($T>0$) phase from the localized ($T<0$) one. The reason is that spatial correlations are absent and the average contribution of the interaction energy vanishes along such a curve. In addition, the MF approximation gives a consistent and semiquantitatively correct description in the whole phase diagram, including the $T=0$ curve and the (metastable) negative $T$ region. This makes our mean-field approximation substantially different from a large-connectivity expansion [22]  [^1] This is because we apply the mean-field approximation exclusively to the mass variables, while treating the phases exactly and keeping nearest-neighbors interactions, see Sec. 2..
In Sec. 2 we present the model and the method. The main results of the MF approximation are discussed making a distinction between positive (Sec. 3) and negative (Sec. 4) temperatures, because in the latter case it is necessary to introduce a cutoff in the mass integral which defines the partition function. All the details of the calculation and of the simulations are deferred to the appendices.
@@ 2 The Model and its Mean Field Theory
We start by defining the DNLS model and recalling the most important properties of its phase diagram. As usual, we consider a one dimensional lattice of size $N$ and associate a local mass $c_{n}>0$ and a phase $\phi_{n}$ to each site $n$. The system Hamiltonian is
@@
In Fig. 1, we plot the numerical values of $h_{nl}$ for the DNLS for different temperatures as a function of the mass density $a$ and for $J=1$. Interestingly, they align almost perfectly along the corresponding isothermal lines of the C2C ($J=0$), showing that given the mass density and the temperature, the nonlinear energy $h_{nl}$ (approximately) coincides in the two models. This implies that the angular degrees of freedom contribute “simply" as an additional interaction energy $h_{int}$, bounded between 2aJ-2aJ in the ground state and 0 on the critical line.
Figure 1: Value of $h_{nl}$ along three isothermal curves ($\beta=0.01,0.1,10$ from top to bottom) comparing the results of equilibrium numerical simulations (see A) of the DNLS equation (symbols) with the prediction of the MF theory (dashed lines) and with the energy of the C2C model (dotted lines) as derived in [16, 23].
Within the grand canonical formalism, the equilibrium properties of the DNLS model can be extracted from the partition function
@@
Figure 2: Relative difference between the MF partition function $z$ computed by numerically integrating the exact expression (9) and its analytical approximation $z_{k}$ obtained as a series expansion at order $k$ in $w$. The various symbols correspond to different $w$ values as from the legend. The value of $\beta$ is always equal to $\pm 0.4$ except for the lowermost curve, where $\beta=-0.01$. Notice that the sign of $\beta$ coincides with the sign of $w$. For negative $\beta$, the numerical integration is performed by imposing the same cutoff $c^{*}$ as in the C2C model.
@@ 3 Positive temperatures
Figure 3: Comparison between MF theory (curves) and exact results obtained via grandcanonical simulations (symbols). Each curve corresponds to a different $T=1/\beta$, increasing from bottom to top according to the legenda. The size of the system is $N=100$ and details about simulations are given in A.
@@
Figure 4: Same as in Fig. 3, but we now compare MF and exact numerical results in the planes $(\mu,a)$ and $(\mu,h)$ rather than in $(a,h)$. In this way a horizontal shift at low temperature is evident, due to a wrong determination of the chemical potential in the MF theory. Symbols: exact results from grandcanonical simulations. Full lines: MF approximation. Dashed lines: MF approximation corrected for the chemical potential: $\mu\to\mu-1$.
@@
Figure 5: The contributions $h_{int}$ and $h_{nl}$ are presented versus the total energy $h$ suitably rescaled (see the main text). Symbols are the exact results from microcanonical simulations, as described in Appendix A, but with the removal of the interaction with the heat bath and starting from an initial condition with the chosen values of $a$ and $h$.
It is worth stressing that the same linear behaviors are expected in the MF approximation, because the limiting values of the two contributions, $y_{nl,int}(0,1)$, are the same, see B and C, and because the exact and MF theories compare very well in the $(a,h)$ plane, see Figs. 1,3. Nevertheless, it is not possible to derive analytically the linear behavior of $y_{nl,int}(x)$ within the MF theory, because such theory works with the grandcanonical variables $\beta,\mu$ and their relation with $a$ and $h$ can be easily obtained numerically, but no closed-form solutions are available. However, we can use the expansions in $w$, which is what we are going to do now.
@@
Figure 6: This figure intends to represent the relationship between the total energy $h$ of the system and its nonlinear part, $h_{nl}$. For small $\beta$ both quantities tend to $h_{c}=2a^{2}$, therefore we plot the relative differences between the two energies and the critical energy. The depedence of this ratio on $m$ is well accounted for by the MF result (dashed line).
@@
It is worth stressing that the approximation $R=4/(4+\pi m^{2}/2)$ is not valid for vanishing $T$. In this limit, $m$ diverges, which would imply a vanishing $R$, while in reality $R$ goes to a constant,
@@ 4 Negative-temperature homogeneous states
Figure 7: Sketch of the potential $U(c)$, Eq. (18), for $\theta=0$ and small, negative $\beta$.
The crossover from the first to the second regime occurs at $c_{1}\simeq\beta^{2}q^{2}/m^{2}$, while the crossover from the second to the third regime occurs at $c_{2}\simeq|m/\beta|$. For the potential to be partially confining, it is therefore necessary that $c_{2}$ is sufficiently larger than $c_{1}$. This implies $|m|\gg|\beta|q^{2/3}$, which, in the limit of small $\beta$, leads to $a\ll 1/|\beta|^{3/4}$ (see Eqs. (11,12) and a comment in the conclusions, next section). In this limit, $|U(c_{1})|\ll U(c_{2})$, so that the anisotropic term is fully negligible and the potential barrier matches the one derived in [10]. In this regime, one can set the cutoff mass equal to the potential maximum, $c^{*}=c_{2}$. However, we also see that for “large" $a$ (and fixed $\beta$), the potential is not locally confining and one must expect breathers to arise much more rapidly. This is at variance with the C2C model, where there is always a minimum in 0.
Figure 8: MF potentials $U(c,\phi=0)=\beta c^{2}+2q\sqrt{c}-mc$, for different values of $\beta$ and $m$. The full (blue) line corresponds to $\beta=0$ and $U(c)$ is linear.
@@ 5 Conclusions
In this paper, we have developed a mean-field approach to derive explicit expressions for the most relevant observables describing the equilibrium state of the DNLS equation. The approximation consists in assuming that the interaction is gauged by the average amplitude of the neighbouring sites rather than by their actual value. It is important that phases are treated exactly.
In a certain sense, the mean-field approximation is a step forward compared to the stochastic C2C model in which the coupling term is completely neglected and phase variables do not appear. In fact, the MF theory is accurate not only in the small $\beta$ region, where the interaction energy is on average negligible: it turns out that the resulting approximation provides a good representation in a much wider region of the parameter space, including the correct determination of the $T=0$ curve,
It is also interesting to notice that the MF model provides an accurate description across the transition from positive to negative temperature, where instabilities are expected to arise. In fact, MF theories typically fail precisely in proximity of phase transitions. Here, the success of MF is due to the fact that the interaction energy is negligible in that area, and the transition itself occurs already in the absence of interactions.
@@
Figure 9: Comparison of exact numerical results (circles) and MF prediction (blue line) for negative $\beta=-0.01$ and different values of m=5,2.5,1.5,1m={-5,-2.5,-1.5,-1}, from left to right. Main: comparison in the plane $(a,h)$. Inset: comparison in the plane $(a,h-2a^{2})$. Here it is manifest we are in the negativeT-T region, because $(h-2a^{2})>0$. We also make evident the statistical error bar for the circles, obtained as spatial and temporal averages of grandcanonical simulations.
Figure 10: As in Fig. 9, but now the comparison is done in the planes $(\mu,a)$ and $(\mu,h)$ rather than in the plane $(a,h)$.
@@ Acknowledgements
SI and PP acknowledge financial support from the Italian MUR PRIN2022 project “Breakdown of ergodicity in classical and quantum many-body systems" (BECQuMB) Grant No. 20222BHC9Z.